On a horizontal axis whose unit is the meter, a linear load ranging from 0 to 1 ma a linear load distribution = 2 nC / m.determine the modulus of the electric field created by the previous loaded bar at the point A of abscissa 2m (we have to find the relation between l, which is the distance between the elementary bar and the point A and x which sweeps the segment [0: 1]

1
Expert's answer
2021-07-23T11:27:01-0400

Electric field

"E=\frac{kq}{r^2}\\\lambda=\frac{q}{l}\\q=\lambda l\\E=\frac{k\lambda l}{r^2}"

Put value

"E=\frac{9\times10^9\times2\times l}{2^2}"

"E=4.5\times10^9\times l" N/c